# ReVision: A Dataset and Baseline VLM for Privacy-Preserving Task-Oriented Visual Instruction Rewriting

### v2

@@ 1 Introduction
The rapid integration of conversational AI into AR, VR, smartphones, and wearables has heightened the demand for multimodal systems that can interpret text, images, speech, and gestures seamlessly. Devices like the Meta Ray-Ban Smart Glasses, Apple Vision Pro, and tools like Google Lens enable users to ask specific questions about their visual surroundings - yet entire images, often containing sensitive background data unrelated to the query, are transmitted to cloud-based large or semi-large vision-language models (VLMs) chen2023pali; liu2023visual; qwen-vl, posing serious privacy risks. This highlights a key challenge: executing task-oriented multimodal commands while preserving user privacy. On-device processing is increasingly favored to avoid exposing private content such as faces, locations, or documents. However, while large VLMs like PaLI-X, LLaVA, and Qwen-VL excel at complex tasks, they are too resource-intensive for local use, and smaller, more private models often lack the broad world knowledge needed for robust multimodal understanding.
To address this, we propose ReVision, an approach based on Visual Instruction Rewriting that converts multimodal instructions into text-only commands. By transforming complex visual interactions into structured text, existing lightweight on-device conversational AI models can efficiently process user instructions without sending either sensitive visual or textual data to external servers. We introduce a curated dataset consisting of $\langle$ image, original instruction, rewritten instruction $\rangle$ triplets, covering diverse real-world tasks. A freshly built, compact, 250-million-parameter vision-language model liu2023visual is fine-tuned on this dataset and evaluated using NLG metrics (such as BLEU, METEOR, ROUGE) and semantic parsing accuracy.
Our findings demonstrate that our compact model achieves an acceptable level of rewriting capabilities, and performs better compared to popular baselines such as PaliGemma-v2 steiner2024paligemma and Qwen2VL wang2024qwen2 in zero-shot settings and a fully fine-tuned version of a 250M parameter VLM (SmolVLM) marafioti2025smolvlm. Additionally, even an 8-bit quantized version of our model (<500MB on storage disk) achieves effective instruction rewrites while maintaining a small computational footprint. We strongly believe this approach bridges the gap between large-scale multimodal AI and privacy-centric, on-device execution, ensuring secure, real-time interaction with AR/VR and smartphone interfaces.
@@
  • A 250M-parameter baseline VLM using the Perceiver Resampler laurenccon269587869matters, pretrained on captioning datasets and fine-tuned on our rewriting dataset.
@@ 2 Related Work
Instruction or query rewriting and semantic parsing have been widely explored in conversational AI to improve query understanding and response generation. Early methods relied on rule-based transformations and supervised learning semantic_parsing_survey, while recent advances leverage LLMs for dynamic query refinement ye2023large; mo2023convgqr. Generative query rewriting frameworks such as LLM-R2 llm-r2 enhance text ranking, and personalized query rewriting methods cho-etal-2021-personalized refine queries based on user preferences. However, these techniques focus primarily on textual query transformations and do not extend to multimodal task-oriented instruction processing. Visual instruction tuning has emerged as a key development in multimodal AI, with models like LLaVA liu2023visual and PaLI-X chen2023pali demonstrating strong vision-language capabilities. While these models excel in multimodal question answering, they are not optimized for rewriting task-oriented instructions. Similarly, Patel et al. patel2020generating explore generating natural questions from images for multimodal assistants, but their work focuses on question generation rather than instruction rewriting. Unlike these approaches, our work introduces a dedicated dataset and a compact model for Visual Instruction Rewriting, specifically designed to convert multimodal user instructions into structured text for privacy-preserving, on-device execution.
The closest work to ours is MARRS ates2023marrs, which integrates multimodal reference resolution with query rewriting to improve conversational grounding. However, MARRS relies on rule-based replacements after reference resolution in a non-VLM setting, whereas our approach focuses on learning-based instruction rewriting to enable structured task execution from multimodal inputs. Other highly relevant studies are by zhang2022can and wei2021visual, which investigate whether open-domain text-based QA systems can handle visual knowledge questions by reformulating them into purely textual queries. Their work highlights the effectiveness of query rewriting in bridging the gap between vision and language using a modular approach different from ours but aligns closely with our goal of rewriting multimodal instructions into structured text. However, while their approach focuses on adapting visual questions for open-domain QA, our work is specifically designed for task-oriented instruction execution, making it applicable to a broader set of real-world multimodal interactions.
@@ 3 Constructing a Dataset for Visual Instruction Rewriting
structured arguments louvan2020recent; aghajanyan2020conversational. For example, when a user says, "Add the team meeting to my calendar for Friday at 3 PM", the system must parse the intent (CreateCalendarEvent) and extract arguments such as the EventTitle (“team meeting”), EventDate (“Friday”), and EventTime (“3 PM”) to schedule the event correctly. Unlike purely text-based interactions, multimodal instructions, particularly those directed at conversational AI assistants on AR/VR devices (e.g., Apple’s Siri for Apple Vision Pro), introduce additional challenges such as ellipsis and coreference resolution. For instance, a user may look at a book cover and ask, “Who wrote this?” or point at a product in an AR interface and say, “How much does this cost?” Traditional text-based semantic parsers struggle with such instructions since critical visual context is missing. Thus, to bridge the gap between multimodal input and existing conversational AI stacks, we introduce a dataset specifically designed for rewriting multimodal instructions into structured text that can be processed by standard text-based semantic parsers. Figure 1 illustrates a representation of the dataset collection requirement, highlighting the transformation of multimodal inputs into text-based rewrites.
To construct our dataset, we first define an ontology of intents and arguments, as existing ontologies in conversational AI and semantic parsing are often proprietary and unavailable for research use. We take inspiration from goel2023presto for ontology and extend it to accommodate multimodal task-oriented interactions. Figure 5 presents an overview of the intents and arguments in our ontology. Next, we curate a diverse set of images covering various real-world multimodal interaction scenarios, including book covers, product packaging, paintings, mobile screenshots, flyers, signboards, and landmarks. These images are sourced from publicly available academic datasets, such as OCR-VQA [^4] https://ocr-vqa.github.io/, CD and book cover datasets, Stanford mobile image datasets [^5] http://web.cs.wpi.edu/~claypool/mmsys-dataset/2011/stanford/, flyer OCR datasets666github.com/Skeletonboi/ocr-nlp-flyer.git, signboard classification datasets777github.com/madrugado/signboard-classification-dataset, Google Landmarks888github.com/cvdfoundation/google-landmark, and Products-10K [^9] https://products-10k.github.io/.
@@
Upon identifying and verifying the images, we employ the GPT-4 model from OpenAI achiam2023gpt to systematically generate and refine multimodal instructions into rewritten text-based instructions. The process begins with a bootstrap phase, where GPT-4 is prompted to generate 20 direct questions per image by explicitly referencing visible objects or textual elements while adhering to the intent list defined in Figure 5. A second prompting phase then validates the generated questions against the corresponding image, filtering out ambiguous or irrelevant instructions to ensure alignment with the visual context.
@@ 4 Developing Small-Scale VLM for Visual Instruction Rewriting
We develop a lightweight vision-language model (shown in Figure 3) tailored for instruction-following tasks by integrating a pretrained vision encoder with an instruction-tuned language model, following the popular multimodal fusion approach zhang2024vision. Since vision encoders and instruction-tuned language models operate in different embedding spaces, a multimodal projector liu2023visual is used to align the encoded image features with the token embedding space of the language model. Our approach is similar to PaLI-Gemma beyer2024paligemma, where an image encoder based on the SigLIP architecture zhai2023sigmoid extracts $D$-dimensional image encodings for $N$ patches from a single input image, say $V_{1},V_{2},...,V_{N}$). Building on laurenccon269587869matters, who demonstrated that using a sampling technique to extract the most relevant $M$ patch encodings from a larger set of $N$ samples improves efficiency, we employ Perceiver Sampler jaegle2021perceiver to downsample the $N$ patch embeddings into $M$ D-dimensional encodings. These image encodings are then mapped into a shared embedding space via a linear multimodal projector, ensuring compatibility with the language model’s $H$-dimensional token embeddings. We fix $K$ at 64. The projected image embeddings ($H_{1},H_{2},...,H_{M}$) are concatenated with the token embeddings extracted from the tokenized textual input ($H_{1},H_{2},...,H_{K}$), where $K$ represents the number of input tokens. The combined embeddings are then processed by the language model to generate responses. To ensure consistency in input representation, we apply image preprocessing, tokenization, and chat template formatting, making the model familiar with structured multimodal input formats.
Although large-scale vision-language models typically involve hundreds of millions of parameters, our focus is on designing a compact and efficient model capable of running on-device. To maintain a parameter budget under 250M, we select a small SigLIP encoder zhai2023sigmoid (google/siglip-base-patch16-256), which processes images of size $256\times 256$ by dividing them into $16\times 16$ patches, with 768 dimensions in hidden layers. The language model is a 150M-parameter instruction-tuned model from OuteAI [^10] https://www.outeai.com (OuteAI/Lite-Mistral-150M-v2-Instruct) based on the Mistral architecture jiang2023mistral, featuring a vocabulary size of 32,768 and a hidden dimension of 768. Since the hidden dimensions of both the vision encoder and the language model are identical, the projector acts purely as a dimensional transformer without altering the shape of the embeddings. While the model’s limited size may impact its ability to handle multi-turn conversations, it is well-suited for single-turn multimodal instruction rewriting tasks. Additionally, since the model is designed for multimodal deixis resolution, it may not be effective for resolving text-only references in extended conversationsates2023marrs.
@@ 4.1 Model Pretraining
To pretrain the model, we adopt an end-to-end training strategy, leveraging datasets from three key sources: (a) LLaVA-ReCap-CC3M, (b) LLaVA-Pretrain, and (c) LLaVA-CC3M-Pretrain-595K. These datasets are curated from large-scale image-text corpora, including LAION schuhmann2021laion400mopendatasetclipfiltered, Conceptual Captions (CC) sharma-etal-2018-conceptual, and SBU NIPS2011_5dd9db5e, which are filtered for balanced concept distribution and enhanced with synthetic captions generated via BLIP to improve vision-language alignment lmms2023llava; liu2023pretrain; liu2023cc3m. Specifically, LLaVA-ReCap-CC3M focuses on re-captioning images to improve concept coverage, while LLaVA-Pretrain consists of 558K image-caption pairs, forming a strong foundational dataset for multimodal alignment. The LLaVA-CC3M-Pretrain-595K dataset, derived from Conceptual Captions 3M, provides a rich set of image-text pairs to enhance model robustness. The total number of examples is thus a little more than 4M. Despite some redundancy in images across datasets, we ensure sufficient data diversity and scale to instill basic image-text alignment capabilities in our pretrained model.
@@ 4.2 Model Fine-Tuning
The motivation for integrating metadata arises from the limitations of small-scale vision-language models (VLMs). Despite being optimized for rewriting tasks, small VLMs struggle with extracting embedded text from images. OCR is a specialized capability distinct from traditional vision-language alignment lamm2024can; nagaonkar2025benchmarking. However, most modern devices are equipped with built-in OCR and image description capabilities, making it practical to supplement the model with external text recognition systems. To systematically evaluate this approach, we present two different metadata extraction:
@@
  • Self_Caption+EasyOCR: We use rewriter models to generate captions themselves using the simple prompt: “Caption this:”. For OCR, we employ EasyOCR [^11] https://github.com/JaidedAI/EasyOCR, a lightweight text extraction model based on the CRAFT algorithm baek2019character, simulating a low-resource on-device setting.
@@
For deployable small VLM baselines, we include Smol-VLM (256M) marafioti2025smolvlm - the smallest publicly available off-the-shelf VLM [^12] https://huggingface.co/HuggingFaceTB/SmolVLM-Instruct to date. We fine-tuned it on our dataset using the same configuration as our primary model, observing steady loss reduction and convergence.
To clarify the distinction between ReVision and a simple captioning + text fusion approach, and to assess the impact of our dataset, we also compare against two TextOnly baselines: (a) Qwen2.5-0.5B qwen2.5, and (b) Mistral-Lite (our custom text backbone), both fine-tuned in a pure text-to-text setting with instructions, EasyOCR outputs, and GPT-4-generated image captions. These comparisons help isolate the benefits of our dataset and design beyond naïve fusion strategies.
@@ 5 Evaluation Metrics
To evaluate our models in Visual Instruction Rewriting, we use standard NLG metrics (BLEU, METEOR, ROUGE) sharma2017nlgeval alongside task-specific semantic parsing evaluations. While NLG metrics assess linguistic similarity, they do not capture functional quality in downstream AI systems. Effective rewriting must also ensure instructions remain interpretable by semantic parsers extracting intent and arguments louvan2020recent. In the absence of an existing parser tailored to our ontology (Figure 5), we employ GPT-4 as a proxy to simulate an on-device parser for intent classification and argument extraction. To evaluate intent and structure preservation, we compare GPT-4-generated parses for both reference and model-generated rewrites. For clarity, we present a collapsed view of intents and arguments. The following metrics are used for evaluation

### v3

@@ 1 Introduction
The rapid integration of conversational AI into AR, VR, smartphones, and wearables has heightened the demand for multimodal systems that can interpret text, images, speech, and gestures seamlessly. Devices like the Meta Ray-Ban Smart Glasses, Apple Vision Pro, and tools like Google Lens enable users to ask specific questions about their visual surroundings - yet entire images, often containing sensitive background data unrelated to the query, are transmitted to cloud-based large or semi-large vision-language models (VLMs) Chen et al. (2024); Liu et al. (2023); Bai et al. (2023), posing serious privacy risks. This highlights a key challenge: executing task-oriented multimodal commands while preserving user privacy. On-device processing is increasingly favored to avoid exposing private content such as faces, locations, or documents. However, while large VLMs like PaLI-X, LLaVA, and Qwen-VL excel at complex tasks, they are too resource-intensive for local use, and smaller, more private models often lack the broad world knowledge needed for robust multimodal understanding.
To address this, we propose ReVision, an approach based on Visual Instruction Rewriting that converts multimodal instructions into text-only commands. By transforming complex visual interactions into structured text, existing lightweight on-device conversational AI models can efficiently process user instructions without sending either sensitive visual or textual data to external servers. We introduce a curated dataset consisting of $\langle$ image, original instruction, rewritten instruction $\rangle$ triplets, covering diverse real-world tasks. A freshly built, compact, 250-million-parameter vision-language model Liu et al. (2023) is fine-tuned on this dataset and evaluated using NLG metrics (such as BLEU, METEOR, ROUGE) and semantic parsing accuracy.
Our findings demonstrate that our compact model achieves an acceptable level of rewriting capabilities, and performs better compared to popular baselines such as PaliGemma-v2 Steiner et al. (2024) and Qwen2VL Wang et al. (2024) in zero-shot settings and a fully fine-tuned version of a 250M parameter VLM (SmolVLM) Marafioti et al. (2025). Additionally, even an 8-bit quantized version of our model (<500MB on storage disk) achieves effective instruction rewrites while maintaining a small computational footprint. We strongly believe this approach bridges the gap between large-scale multimodal AI and privacy-centric, on-device execution, ensuring secure, real-time interaction with AR/VR and smartphone interfaces.
@@
  • A 250M-parameter baseline VLM using the Perceiver Resampler Laurençon et al. (2024), pretrained on captioning datasets and fine-tuned on our rewriting dataset.
@@ 2 Related Work
Instruction or query rewriting and semantic parsing have been widely explored in conversational AI to improve query understanding and response generation. Early methods relied on rule-based transformations and supervised learning Kamath and Das (2019), while recent advances leverage LLMs for dynamic query refinement Ye et al. (2023); Mo et al. (2023). Generative query rewriting frameworks such as LLM-R2 Li et al. (2024) enhance text ranking, and personalized query rewriting methods Cho et al. (2021) refine queries based on user preferences. However, these techniques focus primarily on textual query transformations and do not extend to multimodal task-oriented instruction processing. Visual instruction tuning has emerged as a key development in multimodal AI, with models like LLaVA Liu et al. (2023) and PaLI-X Chen et al. (2024) demonstrating strong vision-language capabilities. While these models excel in multimodal question answering, they are not optimized for rewriting task-oriented instructions. Similarly, Patel et al. Patel et al. (2020) explore generating natural questions from images for multimodal assistants, but their work focuses on question generation rather than instruction rewriting. Unlike these approaches, our work introduces a dedicated dataset and a compact model for Visual Instruction Rewriting, specifically designed to convert multimodal user instructions into structured text for privacy-preserving, on-device execution.
The closest work to ours is MARRS Ates et al. (2023), which integrates multimodal reference resolution with query rewriting to improve conversational grounding. However, MARRS relies on rule-based replacements after reference resolution in a non-VLM setting, whereas our approach focuses on learning-based instruction rewriting to enable structured task execution from multimodal inputs. Other highly relevant studies are by Zhang et al. (2022) and Wei et al. (2021), which investigate whether open-domain text-based QA systems can handle visual knowledge questions by reformulating them into purely textual queries. Their work highlights the effectiveness of query rewriting in bridging the gap between vision and language using a modular approach different from ours but aligns closely with our goal of rewriting multimodal instructions into structured text. However, while their approach focuses on adapting visual questions for open-domain QA, our work is specifically designed for task-oriented instruction execution, making it applicable to a broader set of real-world multimodal interactions.
@@ 3 Constructing a Dataset for Visual Instruction Rewriting
structured arguments Louvan and Magnini (2020); Aghajanyan et al. (2020). For example, when a user says, "Add the team meeting to my calendar for Friday at 3 PM", the system must parse the intent (CreateCalendarEvent) and extract arguments such as the EventTitle (“team meeting”), EventDate (“Friday”), and EventTime (“3 PM”) to schedule the event correctly. Unlike purely text-based interactions, multimodal instructions, particularly those directed at conversational AI assistants on AR/VR devices (e.g., Apple’s Siri for Apple Vision Pro), introduce additional challenges such as ellipsis and coreference resolution. For instance, a user may look at a book cover and ask, “Who wrote this?” or point at a product in an AR interface and say, “How much does this cost?” Traditional text-based semantic parsers struggle with such instructions since critical visual context is missing. Thus, to bridge the gap between multimodal input and existing conversational AI stacks, we introduce a dataset specifically designed for rewriting multimodal instructions into structured text that can be processed by standard text-based semantic parsers. Figure 1 illustrates a representation of the dataset collection requirement, highlighting the transformation of multimodal inputs into text-based rewrites.
To construct our dataset, we first define an ontology of intents and arguments, as existing ontologies in conversational AI and semantic parsing are often proprietary and unavailable for research use. We take inspiration from Goel et al. (2023) for ontology and extend it to accommodate multimodal task-oriented interactions. Figure 5 presents an overview of the intents and arguments in our ontology. Next, we curate a diverse set of images covering various real-world multimodal interaction scenarios, including book covers, product packaging, paintings, mobile screenshots, flyers, signboards, and landmarks. These images are sourced from publicly available academic datasets, such as OCR-VQA [^4] https://ocr-vqa.github.io/, CD and book cover datasets, Stanford mobile image datasets [^5] http://web.cs.wpi.edu/~claypool/mmsys-dataset/2011/stanford/, flyer OCR datasets666github.com/Skeletonboi/ocr-nlp-flyer.git, signboard classification datasets777github.com/madrugado/signboard-classification-dataset, Google Landmarks888github.com/cvdfoundation/google-landmark, and Products-10K [^9] https://products-10k.github.io/.
@@
Upon identifying and verifying the images, we employ the GPT-4 model from OpenAI Achiam et al. (2023) to systematically generate and refine multimodal instructions into rewritten text-based instructions. The process begins with a bootstrap phase, where GPT-4 is prompted to generate 20 direct questions per image by explicitly referencing visible objects or textual elements while adhering to the intent list defined in Figure 5. A second prompting phase then validates the generated questions against the corresponding image, filtering out ambiguous or irrelevant instructions to ensure alignment with the visual context.
@@ 4 Developing Small-Scale VLM for Visual Instruction Rewriting
We develop a lightweight vision-language model (shown in Figure 3) tailored for instruction-following tasks by integrating a pretrained vision encoder with an instruction-tuned language model, following the popular multimodal fusion approach Zhang et al. (2024). Since vision encoders and instruction-tuned language models operate in different embedding spaces, a multimodal projector Liu et al. (2023) is used to align the encoded image features with the token embedding space of the language model. Our approach is similar to PaLI-Gemma Beyer et al. (2024), where an image encoder based on the SigLIP architecture Zhai et al. (2023) extracts $D$-dimensional image encodings for $N$ patches from a single input image, say $V_{1},V_{2},...,V_{N}$). Building on Laurençon et al. (2024), who demonstrated that using a sampling technique to extract the most relevant $M$ patch encodings from a larger set of $N$ samples improves efficiency, we employ Perceiver Sampler Jaegle et al. (2021) to downsample the $N$ patch embeddings into $M$ D-dimensional encodings. These image encodings are then mapped into a shared embedding space via a linear multimodal projector, ensuring compatibility with the language model’s $H$-dimensional token embeddings. We fix $K$ at 64. The projected image embeddings ($H_{1},H_{2},...,H_{M}$) are concatenated with the token embeddings extracted from the tokenized textual input ($H_{1},H_{2},...,H_{K}$), where $K$ represents the number of input tokens. The combined embeddings are then processed by the language model to generate responses. To ensure consistency in input representation, we apply image preprocessing, tokenization, and chat template formatting, making the model familiar with structured multimodal input formats.
Although large-scale vision-language models typically involve hundreds of millions of parameters, our focus is on designing a compact and efficient model capable of running on-device. To maintain a parameter budget under 250M, we select a small SigLIP encoder Zhai et al. (2023) (google/siglip-base-patch16-256), which processes images of size $256\times 256$ by dividing them into $16\times 16$ patches, with 768 dimensions in hidden layers. The language model is a 150M-parameter instruction-tuned model from OuteAI [^10] https://www.outeai.com (OuteAI/Lite-Mistral-150M-v2-Instruct) based on the Mistral architecture Jiang et al. (2023), featuring a vocabulary size of 32,768 and a hidden dimension of 768. Since the hidden dimensions of both the vision encoder and the language model are identical, the projector acts purely as a dimensional transformer without altering the shape of the embeddings. While the model’s limited size may impact its ability to handle multi-turn conversations, it is well-suited for single-turn multimodal instruction rewriting tasks. Additionally, since the model is designed for multimodal deixis resolution, it may not be effective for resolving text-only references in extended conversationsAtes et al. (2023).
@@ 4.1 Model Pretraining
To pretrain the model, we adopt an end-to-end training strategy, leveraging datasets from three key sources: (a) LLaVA-ReCap-CC3M, (b) LLaVA-Pretrain, and (c) LLaVA-CC3M-Pretrain-595K. These datasets are curated from large-scale image-text corpora, including LAION Schuhmann et al. (2021), Conceptual Captions (CC) Sharma et al. (2018), and SBU Ordonez et al. (2011), which are filtered for balanced concept distribution and enhanced with synthetic captions generated via BLIP to improve vision-language alignment Lab (2023); Liu (2023b, a). Specifically, LLaVA-ReCap-CC3M focuses on re-captioning images to improve concept coverage, while LLaVA-Pretrain consists of 558K image-caption pairs, forming a strong foundational dataset for multimodal alignment. The LLaVA-CC3M-Pretrain-595K dataset, derived from Conceptual Captions 3M, provides a rich set of image-text pairs to enhance model robustness. The total number of examples is thus a little more than 4M. Despite some redundancy in images across datasets, we ensure sufficient data diversity and scale to instill basic image-text alignment capabilities in our pretrained model.
@@ 4.2 Model Fine-Tuning
The motivation for integrating metadata arises from the limitations of small-scale vision-language models (VLMs). Despite being optimized for rewriting tasks, small VLMs struggle with extracting embedded text from images. OCR is a specialized capability distinct from traditional vision-language alignment Lamm and Keuper (2024); Nagaonkar et al. (2025). However, most modern devices are equipped with built-in OCR and image description capabilities, making it practical to supplement the model with external text recognition systems. To systematically evaluate this approach, we present two different metadata extraction:
@@
  • Self_Caption+EasyOCR: We use rewriter models to generate captions themselves using the simple prompt: “Caption this:”. For OCR, we employ EasyOCR [^11] https://github.com/JaidedAI/EasyOCR, a lightweight text extraction model based on the CRAFT algorithm Baek et al. (2019), simulating a low-resource on-device setting.
@@
For deployable small VLM baselines, we include Smol-VLM (256M) Marafioti et al. (2025) - the smallest publicly available off-the-shelf VLM [^12] https://huggingface.co/HuggingFaceTB/SmolVLM-Instruct to date. We fine-tuned it on our dataset using the same configuration as our primary model, observing steady loss reduction and convergence.
To clarify the distinction between ReVision and a simple captioning + text fusion approach, and to assess the impact of our dataset, we also compare against two TextOnly baselines: (a) Qwen2.5-0.5B Yang et al. (2024), and (b) Mistral-Lite (our custom text backbone), both fine-tuned in a pure text-to-text setting with instructions, EasyOCR outputs, and GPT-4-generated image captions. These comparisons help isolate the benefits of our dataset and design beyond naïve fusion strategies.
@@ 5 Evaluation Metrics
To evaluate our models in Visual Instruction Rewriting, we use standard NLG metrics (BLEU, METEOR, ROUGE) Sharma et al. (2017) alongside task-specific semantic parsing evaluations. While NLG metrics assess linguistic similarity, they do not capture functional quality in downstream AI systems. Effective rewriting must also ensure instructions remain interpretable by semantic parsers extracting intent and arguments Louvan and Magnini (2020). In the absence of an existing parser tailored to our ontology (Figure 5), we employ GPT-4 as a proxy to simulate an on-device parser for intent classification and argument extraction. To evaluate intent and structure preservation, we compare GPT-4-generated parses for both reference and model-generated rewrites. For clarity, we present a collapsed view of intents and arguments. The following metrics are used for evaluation